OpenStudy (madgirlwithabluebox):

What value completes the square for the expression? x² – 12x

OpenStudy (unklerhaukus):

Some complete square will look like \[(x+a)^2\\~\\\text{if we expand this}\\~\\=(x+a)(x+a)\\=x(x+a)+a(x+a)\\=x^2+ax+ax+a^2\\~\\\text{we get}\\~\\=x^2+2ax+a^2\]

OpenStudy (madgirlwithabluebox):

so how do i start this? im still really confused

OpenStudy (unklerhaukus):

or, for \(b=2a\), ie \(a=\frac b2\) \[(x+\tfrac b2)^2=x^2+bx+(\tfrac{b}{2})^2\]

OpenStudy (unklerhaukus):

you kinda have to use this idea, backwards

OpenStudy (unklerhaukus):

You have \[x^2-12x\] Compare this with \[x^2+bx+(\tfrac b2)^2\]

OpenStudy (unklerhaukus):

The value that completes the square is equal to \((\frac b2)^2\)

OpenStudy (unklerhaukus):

So identify \(b\) from your equation,

OpenStudy (madgirlwithabluebox):

-12

OpenStudy (unklerhaukus):

right, so now compute \((\tfrac b2)^2\)

OpenStudy (madgirlwithabluebox):

36

OpenStudy (unklerhaukus):

good, we now have \[x^2-12x+36\] let's check that it completes the square

OpenStudy (unklerhaukus):

if we break up that middle term in half \[=x^2-6x-6x+36\] and reverse the expansion \[=x(x-6)-6(x-6)\\ =(x-6)(x-6)\\ =(x-6)^2\]

OpenStudy (unklerhaukus):

Does this start to make sense?

OpenStudy (madgirlwithabluebox):

yes!

OpenStudy (unklerhaukus):

if you had an equation to start with, say \[x^2-12x= 100\] Completing the square would show \[(x-6)^2=136\]